# Multilingual Contextualization of Large Language Models for Document-Level Machine Translation

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a wide range of natural language processing tasks (Ouyang et al., 2022; Sanh et al., 2021; Wei et al., 2021; Touvron et al., 2023; Jiang et al., 2023), including machine translation (MT) (Zhang et al., 2023; He et al., 2024; Hendy et al., 2023). Recent work (Alves et al., 2024; Xu et al., 2023) has convincingly shown that LLM-based systems consistently outperform specialized encoder-decoder MT systems for many language pairs (Kocmi et al., 2024; Deutsch et al., 2025). However, the focus of these studies has been primarily on sentence- and paragraph-level translation, overlooking the complexities of translating entire documents where maintaining coherence, consistency, and discourse structure is fundamental. For document-level translation, various techniques have emerged, such as context-aware prompting (Wang et al., 2023a; Wu et al., 2024) and agent-based translation strategies (Wu et al., 2024; Wang et al., 2024). On the other hand, supervised fine-tuning (SFT) has proven highly effective for improving sentence-level MT (Alves et al., 2024; Xu et al., 2023), but its adaptation to document-level translation, and its comparison with other techniques, remain open questions.
@@
    DocBlocks dataset: We release a document-level parallel dataset for tailoring LLMs to document translation, designed to support long-range dependency modeling and discourse coherence.
@@ 2 Multilingual Contextualization of LLMs for Document-level MT
Domain Corpus |D| |S| |W| |W|/|D|
news News Commentary v18.1 125K 4.9M 119M 1K
TED Talks IWSLT2016/2017 33.5K 82.4M 82.4M 2.5K
Parliamentary Europarl v10 44.8K 22.3M 542M 12.1K
Novels GuoFeng 22.6K 1.9M 51.4M 2.3K
BWB 196K 10.3M 450M 2.3K
Figure 2: Statistics of datasets used to create DocBlocks, including domain, corpus name, number of documents $|D|$, sentences $|S|$, words $|W|$, and average document length $|W|/|D|$ as a proxy for discourse complexity (Wang et al., 2023a). We also illustrate the distribution of language pairs across the datasets, highlighting their multilingual composition.
@@
All datasets consist of aligned source and target documents, with symmetric preprocessing to preserve alignment. For datasets containing paragraph-level content with document boundary metadata, we grouped paragraphs into full documents and concatenated them up to a maximum length of $32,768$ tokens. To ensure high-quality data, we applied a rigorous cleaning pipeline to DocBlocks. Low-quality translations were identified using automatic filters, including Bicleaner (Ramírez-Sánchez et al., 2020) (threshold: $0.5$) and CometKiwi-23 (Rei et al., 2023) (threshold: $0.65$). We removed entire documents if over $20\%$ of their sentences fell below these thresholds. We further applied language identification with langid (Lui & Baldwin, 2012) to exclude misaligned language pairs. Additional filters removed documents with fewer than $50$ words, more than $10\%$ identical characters/words/numbers, or a source–target length ratio above $1.3$. We also deduplicated all entries.
@@ 2.2 Document-Level Fine-Tuning with DocBlocks
Following prior work (Zhang et al., 2018; Wu et al., 2024), we adopt a two-step learning strategy, where LLMs are first tuned for sentence-level MT and then adapted for document-level MT tasks. Given the availability of sentence-level instruction-tuned LLMs (Alves et al., 2024; Martins et al., 2024), we focus only on the second step, continuing fine-tuning on document-level data. To ensure high-quality translations, we calculate the loss exclusively on target tokens, excluding prompt tokens (source and instruction tokens), thereby preventing penalties for prompt format adherence.
@@ 2.3 Inference
The chunking method translates a document chunk by chunk, with each chunk comprising a fixed number of sentences, paragraphs, or tokens. It can be applied on its own or enhanced in two complementary ways that address different aspects of the translation process. First, contextual chunking improves coherence by conditioning the translation of each chunk on previously translated source-target chunk pairs. Second, quality-aware chunking focuses on selecting the best translation using Minimum Bayes Risk (MBR) decoding (Kumar & Byrne, 2004; Eikema & Aziz, 2022), guided by quality estimation metrics such as COMET (Rei et al., 2022) and ContextCOMET (Vernikos et al., 2022). These metrics may incorporate contextual information and help choose the final output by maximizing expected utility: $\hat{y}_{mbr}=\arg\max_{y\in\bar{\mathcal{Y}}}\mathbb{E}_{{Y}}\left[u({Y},y)\right]\approx\arg\max_{y\in\bar{\mathcal{Y}}}\frac{1}{|\bar{\mathcal{Y}}|}\sum_{y_{t}\in\bar{\mathcal{Y}}}\mathcal{M}([c_{t},y_{t}],[c,y])$, where $\bar{\mathcal{Y}}$ is a list of candidates sampled by the model and $\mathcal{M}$ is the reference-based quality metric, which can optionally use context, $c$.
@@ 3.1 Experimental Setup
In this work, we employ the closed-source GPT-4o (OpenAI et al., 2023) and the open-source Qwen2.5-72B-Instruct (Bai et al., 2023) as our baseline models. These two state-of-the-art LLMs serve as the primary comparisons for evaluating our approach to adapting sentence-level LLMs to document-level MT. We use greedy decoding for all inference methods, except for quality-aware chunking, where we generate $32$ candidates using $p$-nucleus sampling (Holtzman et al., 2019) with $p=0.6$.
@@
To assess the effectiveness of our method across different base models, we experiment with three instruction-tuned LLMs: TowerInstruct-Mistral-7B (Alves et al., 2024), EuroLLM-9B-Instruct (Martins et al., 2024), and Qwen2.5-7B-Instruct (Bai et al., 2023). While TowerInstruct-Mistral-7B offers a $32768$ context window, it struggles with document-level MT due to its pretraining and instruction tuning being focused on short sequences. EuroLLM-9B-Instruct has a $4096$ context window, limiting its performance in document-level MT. In contrast, Qwen2.5-7B-Instruct performs better with its $32768$ context window (see Tables 2 and 3).
@@ 3.2 Results and Analysis
Table 2: Doc2Doc translation results on the GuoFeng dataset.
@@
[…]
DocMT-LLMs deliver high-quality chunking decoding.
Chunking-based decoding enables efficient document translation by processing text in segments. Larger chunks provide more context, bridging sentence- and document-level translation. DocMT-LLMs consistently outperform sentence-level models on d-BLEU and d-COMET (Figure 3), especially at larger chunk sizes. In contrast, TowerInstruct-Mistral-7B and EuroLLM-9B-Instruct degrade as chunk size increases. Document-level models maintain strong performance and often surpass Qwen2.5-72B-Instruct. DocMT-Qwen2.5-7B-Instruct improves quality across most chunk sizes, despite Qwen2.5-7B-Instruct already supporting document-level MT. These results show that training on DocBlocks enhances both document translation and chunking decoding, making DocMT-LLMs more robust.
@@
Figure 4: Evaluation scores for contextual chunking on GuoFeng using a context window of $3$, consistent with the training setup.
Figure 5: DocMT-TowerInstruct scores on GuoFeng with chunking (C), contextual (CC), and quality-aware (QAC) strategies.
@@
FLORES-200 WMT23 TICO-19
@@
Table 5 shows that DocMT-LLMs fine-tuned with DocBlocks retain strong sentence-level performance, with COMET score differences from their sentence-level counterparts typically within $0.5$ points, indicating no catastrophic forgetting. Qwen2.5-7B-Instruct even improves sentence-level performance through document-level training, enhancing robustness without sacrificing sentence-level quality.
@@
DocMT-LLMs consistently outperform (in quality and inference speed) agent-based methods like TRANSAGENTS (Wu et al., 2024) and DELTA (Wang et al., 2024) in both Doc2Doc and contextual chunking translation tasks (Figure 6). Agent-based methods iteratively refine translations using memory-based mechanisms with dictionaries and using a hierarchical workflow. This multi-step pipeline and inter-agent communication can introduce significant latency, slowing inference. In contrast, training a DocMT-LLM requires $120$ hours on $4$ NVIDIA H100100100100 $80$GB GPUs but runs efficiently on a single NVIDIA A6000600060006000 $48$GB GPU afterward. This initial cost is offset after processing just 42 documents of 400 sentences each. Therefore, both contextual chunking and Doc2Doc approaches with DocMT-LLMs achieve a superior balance among context preservation, speed, scalability, and translation quality.
Figure 6: d-BLEU scores on the Guofeng test set, with all chapter translations concatenated into a single document (Wu et al., 2024). External models include LLAMA-MT (Wang et al., 2024), GPT-4 variants (OpenAI et al., 2023), GOOGLE, HW-TSC (Xie et al., 2023), and agent-based approaches such as TRANSAGENTS and DELTA.
Doc2Doc Chunking Contextual Chunking
@@
DocBlocks curation process impacts learning.
[…]
We conducted an ablation study to assess how varying the proportion of sentence-level data in DocBlocks affects translation performance, while keeping document-level data constant. Results shown in Figure 7 demonstrate that just $10\%$ sentence-level data is sufficient to maintain sentence-level quality, while higher proportions yield diminishing returns and reduce document-level performance. Allocating $90\%$ of training data to document-level pairs provides the best balance, enhancing document-level performance without compromising sentence-level capabilities.
@@ 4 Related Work
Traditional sentence-level MT (Sutskever et al., 2014; Bahdanau et al., 2014; Vaswani et al., 2017) often leads to inconsistencies in terminology, tone, and context in longer texts (Xiong et al., 2018; Hu & Wan, 2023; Wang et al., 2023b). Document-level MT (Maruf et al., 2021) improves coherence by incorporating broader context across sentences and paragraphs. Recent approaches include document embeddings (Macé & Servan, 2019; Huo et al., 2020), multi-encoder architectures (Zhang et al., 2018; Voita et al., 2018; Bawden et al., 2018), and enhanced attention mechanisms for long-range dependencies (Zhang et al., 2020; Miculicich et al., 2018; Wu et al., 2023). While promising, progress is limited by the lack of high-quality document-level parallel data (Liu & Zhang, 2020; Wang et al., 2023b), which hinders training for complex domains like literary translation (Ghazvininejad et al., 2018; Toral & Way, 2018). We address this by curating DocBlocks, a high-quality document-level dataset sourced from public data.
@@
LLMs have recently demonstrated strong performance in MT tasks. Their ability to process long contexts opens new possibilities for modeling discourse structure and inter-sentence relationships. Approaches such as context-aware prompting (Wang et al., 2023a), agent-based modeling (Wu et al., 2024; Wang et al., 2024), and fine-tuning strategies (Wu et al., 2024) have been explored to enhance translation quality. Despite these advances, LLMs continue to struggle with maintaining coherent translations over long documents due to challenges in training data and context modeling. To address these challenges, we fine-tune LLMs using DocBlocks, enhancing their multilingual contextualization for document translation while avoiding the cost of full retraining.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a wide range of natural language processing tasks (Ouyang et al., 2022; Sanh et al., 2021; Wei et al., 2021; Touvron et al., 2023; Jiang et al., 2023), including machine translation (MT) (Zhang et al., 2023; He et al., 2024; Hendy et al., 2023). Recent work (Alves et al., 2024; Xu et al., 2023) has convincingly shown that LLM-based systems consistently outperform specialized encoder-decoder MT systems for many language pairs (Kocmi et al., 2024; Deutsch et al., 2025). However, the focus of these studies has been primarily on sentence- and paragraph-level translation, overlooking the complexities of translating entire documents where maintaining coherence, consistency, and discourse structure is fundamental. For document-level translation, various techniques have emerged, such as context-aware prompting (Wang et al., 2023a; Wu et al., 2024) and agent-based translation strategies (Wu et al., 2024b; Wang et al., 2024). On the other hand, supervised fine-tuning (SFT) has proven highly effective for improving sentence-level MT (Alves et al., 2024; Xu et al., 2023), but its adaptation to document-level translation, and its comparison with other techniques, remain open questions.
@@
    DocBlocks dataset [^1] Available on Hugging Face.: We release a document-level parallel dataset for tailoring LLMs to document translation, designed to support long-range dependency modeling and discourse coherence.
@@ 2 Multilingual Contextualization of LLMs for Document-level MT
Domain Corpus Pre-filtering Post-filtering
|D| |S| |W| |W|/|D| |D| |S| |W| |W|/|D|
news News Commentary v18.1 125.0K 4.9M 119.0M 1.0K 110.0K 4.4M 96.4M 0.9K
TED Talks IWSLT2016/2017 33.5K 3.3M 82.4M 2.5K 29.7K 2.6M 64.0M 2.1K
Parliamentary Europarl v10 44.8K 22.3M 542.0M 12.1K 21.3K 10.5M 264.8M 12.4K
Novels GuoFeng 22.6K 1.9M 51.4M 2.3K 18.2K 1.7M 28.5M 1.6K
BWB 196.0K 10.3M 450.0M 2.3K 62.3K 4.2M 110.3M 1.8K
Figure 2: Statistics of datasets used to create DocBlocks, including domain, corpus name, number of documents $|D|$, sentences $|S|$, words $|W|$, and average document length $|W|/|D|$ as a proxy for discourse complexity (Wang et al., 2023a). We also illustrate the distribution of language pairs across the filtered datasets, highlighting their multilingual composition.
@@
All datasets consist of document-level parallel corpora. We do not enforce strict sentence-to-sentence alignment. Instead, alignment is maintained at the document level, allowing for variation in sentence count and structure – reflecting realistic conditions in document translation. This approach enables the model to rely on broader semantic context rather than rigid sentence mappings. For datasets containing paragraph-level content with document boundary metadata, we grouped paragraphs into full documents and concatenated them up to a maximum length of $32,768$ tokens. To ensure high-quality data, we applied a rigorous cleaning pipeline to DocBlocks. Low-quality translations were identified using automatic filters, including Bicleaner (Ramírez-Sánchez et al., 2020) (threshold: $0.5$) and CometKiwi-23 (Rei et al., 2023) (threshold: $0.65$). We removed entire documents if over $20\%$ of their sentences fell below these thresholds. We further applied language identification with langid (Lui & Baldwin, 2012) to exclude misaligned language pairs. Additional filters removed documents with fewer than $50$ words, more than $10\%$ identical characters/words/numbers, or a source–target length ratio above $1.3$. We also deduplicated all entries.
@@ 2.2 Document-Level Fine-Tuning with DocBlocks
Following prior work (Zhang et al., 2018; Wu et al., 2024), we adopt a two-step learning strategy, where LLMs are first tuned for sentence-level MT and then adapted for document-level MT tasks. Given the availability of sentence-level instruction-tuned LLMs (Alves et al., 2024; Martins et al., 2024), we focus only on the second step, which further fine-tunes these models on document-level translation data. To ensure high-quality translations, we calculate the loss exclusively on target tokens, excluding prompt tokens (source and instruction tokens), thereby preventing penalties for prompt format adherence.
@@ 2.3 Inference
The chunking method translates a document chunk by chunk, with each chunk comprising a fixed number of sentences, paragraphs, or tokens. It can be applied on its own or enhanced in two complementary ways that address different aspects of the translation process. First, contextual chunking improves coherence by conditioning the translation of each chunk on previously translated source-target chunk pairs. Second, quality-aware chunking focuses on selecting the best translation using Minimum Bayes Risk (MBR) decoding (Kumar & Byrne, 2004; Eikema & Aziz, 2022), guided by quality estimation metrics such as COMET (Rei et al., 2022) and ContextCOMET (Vernikos et al., 2022). These metrics may incorporate contextual information and help choose the final output by maximizing expected utility: $\hat{y}_{mbr}=\arg\max_{y\in\bar{\mathcal{Y}}}\mathbb{E}_{{Y}}\left[u({Y},y)\right]\approx\arg\max_{y\in\bar{\mathcal{Y}}}\frac{1}{|\bar{\mathcal{Y}}|}\sum_{y_{t}\in\bar{\mathcal{Y}}}\mathcal{M}([c_{t},y_{t}],[c,y])$, where $\bar{\mathcal{Y}}$ is a list of candidates sampled by the model and $\mathcal{M}$ is the reference-based quality metric, which can optionally use context $c$.
@@ 3.1 Experimental Setup
In this work, we employ the closed-source GPT-4o (OpenAI et al., 2023) and the open-source models Qwen2.5-72B-Instruct (Bai et al., 2023) and Llama-3.3-70B-Instruct (Grattafiori et al., 2024) as our baselines. These three state-of-the-art LLMs serve as the primary comparisons for evaluating our approach to adapting sentence-level LLMs to document-level MT. We use greedy decoding for all inference methods, except for quality-aware chunking, where we generate $32$ candidates using $p$-nucleus sampling (Holtzman et al., 2019) with $p=0.6$.
@@
To assess the effectiveness of our method across different base models, we experiment with three instruction-tuned LLMs: TowerInstruct-Mistral-7B (Alves et al., 2024), EuroLLM-9B-Instruct (Martins et al., 2024), and Qwen2.5-7B-Instruct (Bai et al., 2023). While TowerInstruct-Mistral-7B offers a $32768$ context window, it struggles with document-level MT due to its pretraining and instruction tuning, which are focused on short sequences. EuroLLM-9B-Instruct has a $4096$ context window, limiting its performance in document-level MT. In contrast, Qwen2.5-7B-Instruct performs better with its $32768$ context window (see Tables 2 and 3).
@@ 3.2 Results and Analysis
Table 2: Doc2Doc translation results on the GuoFeng dataset. Best overall values are bolded, and group-specific bests are underlined.
@@
DocMT-LLMs deliver high-quality chunking decoding.
Chunking-based decoding enables efficient document translation by processing text in segments. Larger chunks provide more context, bridging sentence- and document-level translation. DocMT-LLMs consistently outperform sentence-level models on d-BLEU and d-COMET (Figure 3), especially at larger chunk sizes. In contrast, TowerInstruct-Mistral-7B and EuroLLM-9B-Instruct degrade as the chunk size increases. Document-level models maintain strong performance and often surpass Qwen2.5-72B-Instruct. DocMT-Qwen2.5-7B-Instruct improves quality across most chunk sizes, despite Qwen2.5-7B-Instruct already supporting document-level MT. These results show that training on DocBlocks enhances both document translation and chunking decoding, making DocMT-LLMs more robust.
@@
Figure 4: Evaluation scores for contextual chunking on GuoFeng using a context window of 33, consistent with the training setup. Improvements demonstrate CAPT’s effect in enabling DocMT-LLMs to better leverage prior translated chunks during decoding.
Figure 5: DocMT-TowerInstruct-7B performance on GuoFeng using chunking (C), contextual chunking ($C$), and quality-aware chunking (QAC), which combines contextual chunking with paragraph-level quality metrics.
@@
Table 5 shows that DocMT-LLMs fine-tuned with DocBlocks retain strong sentence-level performance, with COMET score differences from their sentence-level counterparts typically within $0.5$ points, indicating no catastrophic forgetting. Qwen2.5-7B-Instruct even improves sentence-level performance through document-level training, enhancing robustness without sacrificing sentence-level quality. Detailed results per language pair are provided in Appendix B.
Models FLORES-200 WMT23 TICO-19
@@
DocMT-LLMs consistently outperform (in quality and inference speed) agent-based methods like TRANSAGENTS (Wu et al., 2024b) and DELTA (Wang et al., 2024) in both Doc2Doc and contextual chunking translation tasks (Figure 6). Agent-based methods iteratively refine translations using memory-based mechanisms with dictionaries and using a hierarchical workflow. This multi-step pipeline and inter-agent communication can introduce significant latency, slowing inference. In contrast, training a DocMT-LLM requires $120$ hours on $4$ NVIDIA H100100 8080GB GPUs but runs efficiently on a single NVIDIA A60006000 4848GB GPU afterward. This initial cost is offset after processing just 42 documents of 400 sentences each. Therefore, both contextual chunking and Doc2Doc approaches with DocMT-LLMs achieve a superior balance among context preservation, speed, scalability, and translation quality.
Figure 6: d-BLEU scores on the Guofeng test set, with chapter translations concatenated into a single document (Wu et al., 2024b). The figure compares document-level translation quality across different paradigms: blue for Doc2Doc, green for Contextual Chunking, orange for the DELTA agent-based approach, and gray for the TRANSAGENTS agent-based approach. In addition to evaluating our base models before and after DocMT training, we include strong external baselines representative of each paradigm – TRANSAGENTS with GPT-4 Turbo (Wu et al., 2024b), DELTA with GPT-4o mini (Wang et al., 2024), and HW-TSC (Xie et al., 2023) for Doc2Doc.
DocBlocks curation process impacts learning.
As shown in Table 6, filtering improves model performance across most datasets. The Multiresolution Document-to-Document (MRD2D) technique provides significant benefits for both document-to-document translation and chunking by introducing additional subdocuments during training, especially for the IWSLT2017 dataset. Context-Aware Prompt Tuning (CAPT) is particularly effective for Contextual Chunking, as it provides context-driven examples that improve performance, particularly for the GuoFeng and IWSLT2017 ($x\rightarrowen$) tasks. Combining filtering with MRD2D and CAPT yields the best results overall, boosting translation quality and contextual understanding across document-level translation tasks. We further examine the impact of balancing the DocBlocks dataset, which is naturally skewed toward English-to-Chinese document pairs due to their greater availability. In this setting, we uniformly sample both from and into English across language directions to ensure an even distribution of translation pairs. Our results indicate that enforcing this balance yields no performance gains. As such, we choose to retain as much data as possible rather than discard samples to achieve balance, prioritizing overall model performance.
Models Doc2Doc Chunking Contextual Chunking
@@
We conducted an ablation study to assess how varying the proportion of sentence-level data in DocBlocks affects translation performance, while keeping document-level data constant. The sentence-level portion is sourced from TowerBlocks (Alves et al., 2024) and uniformly sampled across languages. Results shown in Figure 7 demonstrate that just $10\%$ sentence-level data is sufficient to maintain sentence-level quality, while higher proportions yield diminishing returns and reduce document-level performance. Allocating $90\%$ of training data to document-level pairs provides the best balance, enhancing document-level performance without compromising sentence-level capabilities.
@@ 4 Related Work
Traditional sentence-level MT (Sutskever et al., 2014; Bahdanau et al., 2014; Vaswani et al., 2017) often leads to inconsistencies in terminology, tone, and context in longer texts (Xiong et al., 2018; Hu & Wan, 2023; Wang et al., 2023b). Document-level MT (Maruf et al., 2021) improves coherence by incorporating broader context across sentences and paragraphs. Recent approaches include document embeddings (Macé & Servan, 2019; Huo et al., 2020), multi-encoder architectures (Zhang et al., 2018; Voita et al., 2018; Bawden et al., 2018), and enhanced attention mechanisms for long-range dependencies (Zhang et al., 2020; Miculicich et al., 2018; Wu et al., 2023). While promising, progress is limited by the lack of high-quality document-level parallel data (Liu & Zhang, 2020; Wang et al., 2023b), which hinders training in complex domains such as literary translation (Ghazvininejad et al., 2018; Toral & Way, 2018). Recent efforts have drawn on web crawls (e.g., ParaCrawl (Bañón et al., 2020), BWB (Jiang et al., 2022)), news data (e.g., WMT (Kocmi et al., 2024)), parliamentary records (e.g., Europarl (Koehn, 2005)), public talks (e.g., IWSLT TED task (Cettolo et al., 2017)), literary texts (e.g., GuoFeng (Wang et al., 2023b)), and subtitles (e.g., OpenSubtitles (Lison & Tiedemann, 2016)), as well as large multilingual corpora (e.g., UNPC (Ziemski et al., 2016)). However, many of these suffer from domain constraints, noisy alignments, or unclear document boundaries. We address these issues by curating DocBlocks, a high-quality document-level dataset built from public sources, emphasizing clean alignment, coherent document structure, and careful filtering.
@@
LLMs have recently demonstrated strong performance on MT tasks (Alves et al., 2024). Their capacity to process long contexts opens new possibilities for modeling discourse structure and inter-sentence dependencies, which are essential for document-level translation. To leverage these capabilities, new strategies are being explored. These include context-aware prompting, which uses relevant context directly in the prompt to guide translation (Wang et al., 2023a; Wu et al., 2024), and document-level fine-tuning, which adapts instruction-tuned models to achieve stronger performance (Wu et al., 2024; 2024a; Li et al., 2024). In addition, agent-based approaches such as DELTA (Wang et al., 2024) and TRANSAGENTS (Wu et al., 2024b) further improve consistency by maintaining structured memory across sentences, though they introduce substantial inference overhead. Despite these advances, LLMs still face challenges in maintaining coherence over long documents due to limitations in training data (Wu et al., 2024a) and context modeling (Karpinska & Iyyer, 2023). To address these challenges, we fine-tune LLMs using DocBlocks, enhancing their multilingual contextual understanding for document translation without incurring the high cost of full retraining.